The members of a flying club plan to share equally the cost of a $200,000 airplane. The members want to find five more people to join the club so that the cost per person will decrease by $2000. How many members are currently in the club.
Answer by mananth(16946)

The members of a flying club plan to share equally the cost of a $200,000 airplane. The members want to find five more people to join the club so that the cost per person will decrease by $2000. How many members are currently in the club.
...
let the number of members be x
cost of plane = $200,000
cost per person = 200,000/x
..
if 5 are added
x+5 members
cost per person will be 200,000/(x+5)
...
200,000/x - 200,000/(x+5)=2000
x(x+5) LCD
200,000(x+5)-200,000x/(x(x+5)=2000
200,000x+1000000- 200,000x=2000x(x+5)
+1000000=2000x^2+10000x
/2000
500=x^2+5x
x^2+5x-500=0
x^2+25x-20x-500=0
x(x+25)-20(x+25)=0
(x+25)(x-20)=0
x= 20 members
...
m.ananth@hotmail.ca